# Fast and Private Inference of Deep Neural Networks by Co-designing Activation Functions

### v1

@@ 6.5 Scaling to ImageNet
In this section, we evaluate the scalability of our approach on the ImageNet dataset using a ResNet-50 architecture with 23 million parameters. This architecture was previously too large for training with polynomial activation functions [12]. We compare our approach to COINN and CryptGPUand exclude GForce as they do not consider ImageNet.
@@
We plot the inference time in Figure 10. We observe a significant reduction over COINN across all network delays with a 28×28$\times$28 × reduction in the LAN (0.25 ms) and a 90×90$\times$90 × reduction in the WAN (100 ms). Compared to CryptGPU we find that PILLAR + HoneyBadger is the fastest in all network delays by 3×3$\times$3 × on average. PILLAR + ESPN is slightly slower in the LAN, but once again outperforms CryptGPU in the WAN.
@@
We present a summary of the accuracies in Table 5. We observe a much higher encrypted accuracy for PILLAR compared to COINN and thus, our solution is Pareto dominant. For CryptGPU, we use a pre-trained PyTorch model with state-of-the-art accuracy. Therefore, as expected, CryptGPU has an accuracy $3\%$ higher than the model we trained from scratch. We note that with a higher degree polynomial, we were able to train a $79.2\%$ polynomial model. However, this model is not possible to infer in the $64$-bit field used by CrypTen (as higher degrees need more precision by equation 5). We discuss future directions to further improve this result in Section 7.
@@ 7 Discussion
Our experimental evaluation in Section 6 showed our algorithms significantly outperform all related work in inference time when the network latency is high. While state-of-the-art compared to other polynomial training approaches, PILLAR still incurs a minor accuracy degradation compared to standard models with ReLUs. We posit a few directions for future work to further close this gap between polynomials and ReLUs.
@@
By using CrypTen as our backend, we were limited to a $64$-bit ring for cryptographic operations. As discussed in Section 4, this precision determines the degree and range of polynomials we can use. Interesting future work is to increase this precision to enable higher-degree polynomials and study the performance-accuracy trade-off. Our initial results on ImageNet show that we can train up to a degree eight polynomial without suffering escaping activations. However, we could not increase the ring size to study the effect of higher degrees on inference time.
@@ 8 Related Work
Our work focuses on co-designing the activation functions with cryptography by using polynomials. However, the problem of computing polynomials in MPC is of independent interest and has also been studied in the literature. The state-of-the-art in this space is HoneyBadger, as discussed in Section 4. Other notable works include the initial inspiration for HoneyBadger from Damgård et al. [8]. This early approach conducts exponentiation by blinding and reconstructing the number to be exponentiated so the powers can be computed in plaintext [8]. Building off this idea, Polymath constructs a constant round protocol for evaluating polynomials focused on matrices [25]. However, HoneyBadger outperforms Polymath by reducing both the rounds and the number of reconstructions to one.
@@ Appendix A Evaluation of BatchNorm with Polynomials
Appendix B Quantization Aware Polynomial Fitting
@@
Appendix C Datasets and Architectures
Datasets
We use a collection of common benchmark datasets used in the related work we compare to [17, 29]. Each dataset represents an image classification task of varying difficulty. We use the PyTorch version of each dataset with the pre-determined splits into training and test sets.
  • CIFAR-10/100 [21] is a collection of $60,000$ images of size $32\times 32$ pixels. CIFAR-10 contains $10$ classes, each with $6,000$ images (CIFAR-100 has $100$ classes each with $600$ images). In total, this gives $50,000$ training images and $10,000$ test images.
  • ImageNet [10] contains $1.4$ million images of size $256\times 256$ pixels. There are $1000$ different classes, each with around $1000$ images on average. The training set has $1.2$ million images, and the testing set contains $100,000$ images.
Architectures
We choose the following architectures to facilitate comparisons COINN and GForce [17, 29]. This covers models of depth $7$ to $110$ layers with the number of trainable parameters ranging from $0.2$ to $23$ million.
  • MiniONN [24] is a smaller convolutional model tailored for secure inference. It contains six convolutional layers and one fully connected layer.
  • VGG-16 [31] is a classic deep convolutional network commonly evaluated in the literature [30]. We use the version evaluated in GForce [29], which changes the fully-connected layer from 4096 to 512 neurons.
  • ResNet-18, 32, 50, 110 [16]. ResNets are convolutional neural networks that introduce skip connections between layers that connect non-adjacent layers. They are the most accurate of all the models we consider. We use various sizes following related work [17].

### v2

@@ 6.5 Scaling to ImageNet
In this section, we evaluate the scalability of our approach on the ImageNet dataset using a ResNet-50 architecture with 23 million parameters. This architecture was previously too large for training with polynomial activation functions [12]. We compare our approach to Cheetah, COINN and CryptGPUand exclude GForce as they do not consider ImageNet.
@@
We plot the inference time in Figure 10. We observe a significant reduction over Cheetah and COINN across all network delays. We outperform Cheetah by 39×39$\times$39 × in the LAN (0.25 ms) and 15×15$\times$15 × in the WAN (100 ms). Over COINN, we observe a 28×28$\times$28 × reduction in the LAN and a 90×90$\times$90 × reduction in the WAN. Compared to CryptGPU we find that PILLAR + HoneyBadger is the fastest in all network delays by 3×3$\times$3 × on average. PILLAR + ESPN is slightly slower in the LAN, but once again outperforms CryptGPU in the WAN.
@@
We present a summary of the accuracies in Table 5. We note that Cheetah [17] did not measure accuracy in either their code or paper so we omit them from this comparison. We observe a much higher encrypted accuracy for PILLAR compared to COINN and thus, our solution is Pareto dominant. For CryptGPU, we use a pre-trained PyTorch model with state-of-the-art accuracy. Therefore, as expected, CryptGPU has an accuracy $3\%$ higher than the model we trained from scratch. We note that with a higher degree polynomial, we were able to train a $79.2\%$ polynomial model. However, this model is not possible to infer in the $64$-bit field used by CrypTen (as higher degrees need more precision by equation 10). We discuss future directions to further improve this result in Section 7.
@@
6.6 Real WAN Evaluation
Dataset / Model ESPN HoneyBadger CryptGPU CIFAR-10 / ResNet-110 49.3 $\pm$ 0.1 49.0 $\pm$ 0.1 242.5 $\pm$ 0.9 CIFAR-10 / ResNet-18 15.3 $\pm$ 0.2 12.9 $\pm$ 0.1 48.9 $\pm$ 0.1 CIFAR-100 / ResNet-18 15.4 $\pm$ 0.2 12.9 $\pm$ 0.1 48.9 $\pm$ 0.5 CIFAR-100 / ResNet-32 14.1 $\pm$ 0.1 14.0 $\pm$ 0.1 75.4 $\pm$ 1.6 ImageNet / ResNet-50 153.9 $\pm$ 1.0 104.9 $\pm$ 0.8 268.67 $\pm$ 1.3
Table 6: Real World WAN evaluation. We report total inference time in seconds.
This section gives benchmarks for ESPN, HoneyBadger, and CryptGPU [35] in a real WAN. We use two AWS EC2 g4dn.metal instances, one in the Ohio data centre and one in Frankfurt, Germany. Each machine has $96$ cores, $384$ GB of memory, $100$ GB/s network bandwidth, and a NVIDIA T4 GPU. In practice, we measured 10.8MB/s bandwidth between the two instances. We run all the ResNet architectures and summarize the results in Table 6. We repeat each experiment 20 times and report the mean and 95% confidence intervals. We observe similar trends to the simulated WAN used in previous experiments. Namely, ESPN and HoneyBadger perform similarly in runtime for most models. The exception is larger models like ResNet50, where the bandwidth limits impact ESPN more than HoneyBadger. In all cases, CryptGPU is significantly outperformed by both approaches.
@@ 7 Discussion
Our experimental evaluation in Section 6 showed our algorithms significantly outperform all related work in WAN inference time. While state-of-the-art compared to other polynomial training approaches, PILLAR still incurs a minor accuracy degradation compared to standard models with ReLUs. We posit a few directions for future work to further close this gap between polynomials and ReLUs.
@@
By using CrypTen as our backend, we were limited to a $64$-bit ring for cryptographic operations. As discussed in Section 4.3, this precision determines the degree and range of polynomials we can use (due to either the failure probability of truncation or severe truncation of intermediate values). Interesting future work is to increase this precision to enable higher-degree polynomials and study the performance-accuracy trade-off. Our initial results on ImageNet show that we can train up to a degree eight polynomial without suffering escaping activations. However, we could not increase the ring size to study the effect of higher degrees on inference time.
@@ 8 Related Work
Our work focuses on co-designing the activation functions with cryptography by using polynomials. However, the problem of computing polynomials in MPC is of independent interest and has also been studied in the literature. The state-of-the-art in this space is HoneyBadger, as discussed in Section 4. Other notable works include the initial inspiration for HoneyBadger from Damgrd et al. [8]. This early approach conducts exponentiation by blinding and reconstructing the number to be exponentiated so the powers can be computed in plaintext [8]. Building off this idea, Polymath constructs a constant round protocol for evaluating polynomials focused on matrices [27]. However, HoneyBadger outperforms Polymath by reducing both the rounds and the number of reconstructions to one.
@@ Appendix A Evaluation of BatchNorm with Polynomials
Appendix B Evaluation of Sigmoid with Polynomials
ReLU Sigmoid
Standard Activation 94.7 $\pm$ 0.08 90.2 $\pm$ 0.11
Polynomial Approx. 93.4 $\pm$ 0.16 85.9 $\pm$ 0.11
Table 8: Comparing the effect of the activation function on a ResNet18 model.
In this work, we focus on the ReLU activation function, the default in common architectures such as ResNets [16]. Another reason we focus on ReLU is that it gives better accuracy than alternatives like Sigmoid. In this section, we highlight this accuracy advantage by the accuracy of a ResNet18 model with different activation functions. In Table 8, we evaluate both ReLU and Sigmoid with and without using polynomial approximation. In all cases, we use the ResNet18 architecture with default parameters given in Section 6.1, we note the polynomial is of degree $n=4$. The results are averaged over five random seeds and shown with $95\%$ confidence intervals. We find that ReLU consistently outperforms Sigmoid with and without using polynomial evaluations. However, the accuracy of Sigmoid decreases more, relative to ReLU, when using polynomial approximation. This result further motivates our use of ReLU. We leave further investigation of Sigmoid and other activations for future work.
Appendix C Quantization Aware Polynomial Fitting
@@
Polynomial Input Distribution After Regularization
We plot the histogram of the input to all activation functions of a ResNet 18 model on CIFAR-10 in Figure 12.
Figure 12: Polynomial Input Distribution of ResNet18 on CIFAR10.
Appendix D Communication and Rounds Benchmark
Dataset Model ESPN HoneyBadger CryptGPU COINN GForce Cheetah
CIFAR-10 ResNet-18 0.46 (38) 0.23 (38) 0.45 (174) / / /
ResNet-110 0.55 (221) 0.13 (221) 0.53 (1093) 6.8 / /
VGG-16 0.32 (31) 0.26 (31) / / 0.050 /
CIFAR-100 ResNet-18 0.46 (38) 0.23 (38) 0.45 (174) / / /
ResNet-32 0.16 (65) 0.037 (65) 0.15 (313) 1.9 / /
VGG-16 0.32 (31) 0.26 (31) / / 0.050 /
ImageNet ResNet-50 7.85 (160) 4.04 (160) 7.70 (552) 122.0 / 2.36 (1042)
Table 9: Evaluation of communication in GB and the number of rounds (shown in parenthesis).
We study two additional evaluation metrics of rounds and communication in this section. Recall that the number of rounds significantly affects the protocol latency over WAN. The communication affects each round’s throughput, depending on the network bandwidth. In Table 9, we summarize the communication in GB and the number of rounds (shown in parenthesis) for our work and the related work we compare to in Section 6. We note that both COINN and GForce do not log the communication or rounds in their code base. Thus, we report their communication numbers from the corresponding tables in the papers (COINN [18, Table 3] and GForce [31, Table 7]). Neither work evaluates the number of rounds.
In all cases, ESPN and HoneyBadger significantly dominate all evaluated related work in the number of rounds with a 35×3-5$\times$3 - 5 × improvement. The communication of ESPN and HoneyBadger is significantly less than COINN, approximately the same as CryptGPU, and more than GForce and Cheetah. Our HoneyBadger solution has approximately 2×2$\times$2 × the communication of Cheetah, but Cheetah has 10×10$\times$10 × the rounds; thus, in practice, HoneyBadger gives much better performance as shown in Section 6. GForce has an impressively low communication online (50MB) due to offloading 20GB of communication to an offline phase. Once again, despite having higher communication, we recall that both ESPN and HoneyBadger dominate related work in runtime in the WAN (as shown in Section 6).
Appendix E Correctness Proofs
We begin by proving the correctness of Algorithm 1.
Theorem E.1.
Given an input $[[x]]$ = $x_{A}+x_{B}$ and exponent $k$, Algorithm 1, correctly returns $[[x^{k}]]$.
Proof.
We begin with party $A$. In line 4, they compute their share of the vector $\mathbf{\mathbf{a}}$ where $\mathbf{a}_{i}=x_{A}^{k-i}$ (since $\mathbf{\mathbf{a}}$ was initialized to zero and the for loop iterates over each entry of $\mathbf{\mathbf{a}}$ exactly once). Similarly, in line 5, party $B$ computes their share of $\mathbf{\mathbf{b}}$ where $\mathbf{b}_{i}={k\choose i}x_{B}^{i}$. We recall that party $B$’s share of $\mathbf{a}$ is the zero vector, and similarly for party $A$’s share of $\mathbf{b}$ Then, the vector $\mathbf{p}$ is obtained by multiplying $\mathbf{a}$ and $\mathbf{b}$ element wise in line 6. Thus, $\mathbf{p}_{i}=\mathbf{a}_{i}\cdot\mathbf{b}_{i}=x_{A}^{k-i}{k\choose i}x_{B}^{i}$. The final step (line 7), simply sums $\mathbf{p}$. Therefore,
$s=\sum_{i=1}^{k}\mathbf{p_{i}}=\sum_{i=1}^{k}{k\choose i}x_{A}^{k-i}x_{B}^{i}$ (8)
which applying the binomial theorem (2) gives $(x_{A}+x_{B})^{k}=[[x^{k}]]$. ∎
Given the correctness of Algorithm 1, we now prove the correctness of Algorithm 2. To begin, we bound the failure probability of each truncation step in Algorithm 2.
Theorem E.2.
Consider computing a degree $n$ polynomial fitted to the range $[-\lambda,\lambda)$. Let the global precision (size of the ring) be $L$-bit and the working precision of each value be $p$-bit. Then, the truncation in line 5 of Algorithm 2, fails (The local division of $[[x]_{A},[x]_{B}]*2^{-\bar{s}}\neq[[x*2^{-\bar{s}}]]$) with probability at most
[…]
Proof.
Consider the first truncation of Algorithm 2 in line 5. The input to this truncation is $x$ which we assume is contained in the range $[-\lambda,\lambda)$ with a working precision of $p$-bits. Therefore, the size of $x$ is $2^{\lceil\log_{2}{\lambda}\rceil+1}$ in the integer part and $2^{p}$ in the decimal part. Which gives, $|x|\leq 2^{\lceil\log_{2}{\lambda}\rceil+1+p}$. Given that we are working in a $L$-bit ring and the probability of failure of the truncation protocol is bounded by $|x|/Q$ where $Q$ is the ring size [22], the result follows. ∎
Theorem E.3.
Consider computing a degree $n$ polynomial fitted to the range $[-\lambda,\lambda)$. Let the global precision (size of the ring) be $L$-bit and the working precision of each value be $p$-bit. Then, the truncation in line 7 of Algorithm 2 fails (The local division of $[[x]_{A},[x]_{B}]*2^{-p*(i-1)+\bar{s}\cdot i}\neq[[x*2^{-p*(i-1)+\bar{s}\cdot i}]]$) with probability at most
[…]
where $i$ is the power of $x$ being truncated ($i\leq n$).
Proof.
Consider the truncation in line 7 of Algorithm 2. The input to this truncation is the output of the previous truncation in line 5, raised to the power $i$. We assume the previous truncation was correct. Then after the truncation,
$|x^{\prime}_{i}|=\frac{2^{\lceil\log_{2}{\lambda}\rceil+1+p}}{2^{\lceil(i-2)p/i\rceil}}\leq\frac{2^{\lceil\log_{2}{\lambda}\rceil+1+p}}{2^{(i-2)p/i}}$ (11)
for $i\in\{2,\dots,n\}$, where the inequality holds because $(i-2)p/i$ is positive. After applying the exponentiation by $i$ we get $|\mathbf{p}_{i}|\leq 2^{i(\lceil\log_{2}{\lambda}\rceil+1)+2p}$. Given that we are working in a $L$-bit ring and the probability of failure of the truncation protocol is bounded by $|x|/Q$ where $Q$ is the ring size [22], the result follows. ∎
Theorem E.4.
Given an input $[[x]]$ = $x_{A}+x_{B}$ and polynomial coefficients $\alpha$, Algorithm 2, correctly returns the polynomial evaluation $[[\sum_{i=0}^{n}\alpha_{i}\cdot x^{i}]]$ except with probability.
[…]
Proof.
First, prove correctness assuming the truncation operators are correct (then we will account for the probability of failure). We begin by proving that $\mathbf{p^{\prime}_{i}}=x^{i}$ for $i\in\{1,\dots,n\}$ after line 7. We note that $x$ is actually $x*2^{p}$ due to the fixed point encoding. $\mathbf{p_{1}}$ follows trivially. For $i\in\{2,\dots,n\}$, we work backwards, from line 7 to line 4 expanding the definition of $\mathbf{p^{\prime}_{i}}$
$\mathbf{p^{\prime}_{i}}$ $=$ $\mathbf{p}_{i}*2^{-p*(i-1)+\bar{s}\cdot i}$ (12)
$=$ $(x^{\prime}_{i})^{i}*2^{-p*(i-1)+\bar{s}\cdot i}$ (13)
$=$ $(x\cdot 2^{p}\cdot 2^{-\bar{s}})^{i}*2^{-p*(i-1)+\bar{s}\cdot i}$ (14)
$=$ $x^{i}\cdot 2^{p}$ (15)
where the first lines, up to (14), come from the substitution of lines 7 through 5, respectively. The remaining step follows from basic algebra.
Finally, we bound the failure probability of the algorithm. The truncation steps in line 5 and line 7 each introduce a possibility of wrap around error. Each truncation is executed $n-1$ times. Thus, applying the bounds derived in Theorem E.2 and Theorem E.3 for each $i$ in the for loop, the total failure probability follows. ∎